Jack is now four times as old as his dog. In six years he'll be only twice as old as his dog. How old is Jack? How old is his dog?
Answer by dfrazzetto(283)


J = 4D
J + 6 = 2(D+6)
substituting, J = 4D:
4D + 6 = 2D + 12
2D = 6
D = 3
J = 12
CHECK:
12 = 4(3)
12 + 6 = 2(3 + 6)
18 = 2(9) = 18